OpenStudy (anonymous):

I need help for x= 3y+6

OpenStudy (anonymous):

no

OpenStudy (anonymous):

in the form y=mx+b , m is the slope of the line, whose equation you currently have written as x-3y=6 b is the y intercept of the line, so if the intercept (b) was say 6 , then at the point (0,6) your line would cross the y-axis but you MUST have it in the form "y=mx+b" in order for any of this to work, so, solve for y: x-3y=6 add 3y to both sides x=3y+6 subtract 6 from both sides x-6=3y divide by 3 to solve for y y=(x-6)/3 y=(1/3)x - 2 so m=1/3 which is your slope and b=-2 so that's the y coordinate of where your line cross the y axis

OpenStudy (anonymous):

he is right

OpenStudy (anonymous):

b=-2